OpenStudy (anonymous):

Solve. 2x . 3=5x-5/7

OpenStudy (anonymous):

is the . a multiplication sign?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if it is then you get 6x=5x-(5/7)

OpenStudy (anonymous):

and subtract 5x from both sides, and you get... (I am leaving the answer open so you can figure it out)

OpenStudy (anonymous):

6x-5x=-5/7 x=-5/7